OpenStudy (anonymous):

what would be the equation -8,0 -2,3 0,4

OpenStudy (anonymous):

Do you mean, "Find the equation of the line passin through the points (-8, 0), (-2, 3), and (0, 4)?

OpenStudy (anonymous):

If so, then you should pick two points and use the two-point formula.

OpenStudy (anonymous):

yes that's what is it. my question is saying what wouls be the equation for line h and the points are -8,0 -2,3 and 0,4

OpenStudy (anonymous):

That is, First find the slope.

OpenStudy (anonymous):

I don't know the slope

OpenStudy (anonymous):

The formula for slope is \[(y1-y2)/(x1-x2)\]

OpenStudy (anonymous):

x1, y1 is the first point, and x2, y2 is the second.

OpenStudy (anonymous):

okay can you help me wit that please.

OpenStudy (anonymous):

First, pick two points.

OpenStudy (anonymous):

It doesn't matter which.

OpenStudy (anonymous):

okay im picking point -8,0 and -2,3 what do I di from there

OpenStudy (anonymous):

Then take the first y-coordinate and subtract from it the second.

OpenStudy (anonymous):

Take the first x-coordinate and subtract from it the second.

OpenStudy (anonymous):

okay I got -3 is that correct?

OpenStudy (anonymous):

because 0-3= -3

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

so what do I do from there

OpenStudy (anonymous):

Then do the same with the x's.

OpenStudy (anonymous):

okay so will it be -6

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

okay so what do I do from here. how do I find the equation

OpenStudy (anonymous):

Now divide the difference in the y's by the difference in the x's.

OpenStudy (anonymous):

That is your slope.

OpenStudy (anonymous):

okay I got 2. so how do I find the equation of the coorindinates

OpenStudy (anonymous):

The equation is y-y1=m(x-x1).

OpenStudy (anonymous):

y1 and x1 are numbers, not variables.

OpenStudy (anonymous):

im confused now because all of my other equations are like y=x+2 or something like that

OpenStudy (anonymous):

You can turn it into that, too.

OpenStudy (anonymous):

I just need to find the equation of the lines and the points on the grid

OpenStudy (anonymous):

Using algebraic simplification, you can turn it into standard form. What are you using the equation for?

OpenStudy (anonymous):

well I have a 4 quadrant graph and theres a line with a increasing slope going from quadrant 2 to 1. the plots are the ones I gave you and the question is asking what whould be the eaquation for line h.

OpenStudy (anonymous):

Is the question specifically asking for the equation to be given in standard form?

OpenStudy (anonymous):

Is this an online thing?

OpenStudy (anonymous):

i don't know im only in 7th grade it just says what wuld be the equation. all my other equations are like y=x+2 or times 2. no its my math homework

OpenStudy (anonymous):

I think that the teacher would probably accept the equation that way, but if you're worried, I can show you how to put it into standard form.

OpenStudy (anonymous):

I am in seventh grade, too!

OpenStudy (anonymous):

YES PLEASE!!!!!!!!

OpenStudy (anonymous):

First, you have to multiply all of the (x-x1) stuff by the m stuff. (the slope is m.)

OpenStudy (anonymous):

how come theres no specific equation like y= x+3 or spmething lke that cause that's how all my others are like

OpenStudy (anonymous):

First thing first, get your equation into shape. Replace the x1 and y1 with numbers!

OpenStudy (anonymous):

Besides, the equation that I gave you IS specific!

OpenStudy (anonymous):

okay so how would i do that

OpenStudy (anonymous):

Just because the slope isn't 1 doesn't mean that it isn't specific.

OpenStudy (anonymous):

First things first, get your equation up. Type it in without x1's or y1's.

OpenStudy (anonymous):

sorry i have to go to bed now thanks for your help

OpenStudy (anonymous):

Okay.